OpenStudy (anonymous):

The sum of two consecutive integers is 203. What are the integers?

OpenStudy (anonymous):

Let one integer equal x. The next integer (since they are consecutive) will by x+1. x+(x+1)=203 2x+1=203 2x=202 x=101 So if the first integer is 101, then the next integer will be 102. Then, test to see if values are correct. 101+102=203.

OpenStudy (anonymous):

x = first integer x + 1 = 2nd integer x + x + 1 = 203 -- combine like terms 2x + 1 = 203 -- subtract 1 from both sides 2x = 203 - 1 2x = 202 x = 202/2 x = 101 x + 1 = 101 + 1 = 102 answer : first integer = 101 and 2nd integer = 102 101 + 102 = 203 203 = 203 (correct)

OpenStudy (anonymous):

de ja vous....you got the same thing I got...even the explaining part. Great minds think alike...lol

OpenStudy (anonymous):

you just type faster...lol